OpenStudy (anonymous):

(3^2)^x - 3^x+1 = -1 (need help, step wise) to solve

OpenStudy (anonymous):

using logs

OpenStudy (anonymous):

use pythagoras therem

OpenStudy (anonymous):

dats sum fancy crap rite there

terenzreignz (terenzreignz):

\[\Large \left(3^2\right)^x- 3^{x+1}=-1\] Daunting, isn't it? :D How about this: \[\Large (3^{x})^2- 3^{x+1}=-1\] Less daunting?

terenzreignz (terenzreignz):

@LALFAN4E Ever heard of quadratic equations? This isn't one, obviously, but it certainly *looks* like one, wouldn't you say? ^_^

OpenStudy (anonymous):

@terenzreignz was on the right path, keeping the 1 on the left side though, you expand and then use the quadratic equation once simplified, but thanks guys